What volume of a 1.420 M NaOH solution is required to titrate 25.00 mL of a 4.50 M H,SO, solution?


1
Expert's answer
2020-12-21T03:55:23-0500

Solution:

"Chemical\; reaction:"

"H_2SO_4(aq)\;+\;2NaOH(aq)\implies Na_2SO_4(aq)\;+\;H_2O(l)"


1) By using the acid-base equation:


"M_1V_1=M_2V_2"


"4.5M*25ml=1.42M*V_2"


"V_2=\frac{(4.5M*25ml)}{1.42M}"


Remember 1 mole H2SO4 is equivalent to 2 moles NaOH, that is Normality of H2SO4 = (Molarity x 2)


"79.2*2ml=158ml\;of\;1.42M\;NaOH\;will\;be\;required."


Answer:

"V_2=158.4ml"